40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log, Page 103 of 344  Not logged in ELOG logo
ID Date Author Type Category Subject
  12148   Fri Jun 3 13:05:18 2016 ericqUpdateCDSCDS Notes

Some CDS related things:


Keith Thorne has told us about a potential fix for our framebuilder woes. Jamie is going to be at the 40m next week to implement this, which could interfere with normal interferometer operation - so plan accordingly. 


I spent a little time doing some plumbing in the realtime models for Varun's audio processing work. Specifically, I tried to spin up a new model (C1DAF), running on the c1lsc machine. This included:

  • Removing the unused TT3 and TT4 parts from the IOO block in c1ass.mdl, freeing up some DAC outputs on the LSC rack
  • Adding an output row to the LSC input matrix which pipes to a shared memory IPC block. (This seemed like the simplest way for the DAFI model to have access to lots of signals with minimal overhead).
  • Removing two unused ADC inputs from c1lsc.mdl (that went to things like PD_XXX), to give c1daf.mdl the required two ADC inputs - and to give us the option of feeding in some analog signals.
  • Editing the rtsystab file to include c1daf in the list of models that run on c1lsc
  • Editing the existing DAFI .mdl file (which just looked like an old recolored cut-n-paste of c1ioo.mdl) to accept the IPC and ADC connections, and one DAC output that would go to the fibox. 

The simple DAFI model compiled and installed without complaint, but doesn't succesfully start. For some reason, the frontend never takes the CPU offline. Jamie will help with this next week. Since things aren't working, these changes have not been commited to the userapps svn. 

  12147   Fri Jun 3 12:53:44 2016 ericqUpdateElectronicsCommon board Op amp input offsets

I replaced some of the AD829s with other AD829s, but the offset situation didn't improve.

However, I figured that we don't really need the ~100MHZ bandwidth of the AD829, since the IMC loop limits us to a ~10kHz CARM bandwidth. Also, since we don't routinely use IN2 for anything, I felt free to try something else. 

Specifically, I replaced all of the positive gain AD829s in the input 2 gain ladder with OP27s (U8B->U12B on D1500308), which should have input offset voltages ~30x lower than the AD829s. 

Here is a comparison of the outputs these configurations perform, normalized to the output at the +0dB gain setting - where all of the op amps in the gain ladder are bypassed. 

So, most of the transitions now result in an output offset change of less than 0.5mV, which is nice.

The exception seems to be where the +8dB stage is switched in or out. I may try replacing this one, as these transitions cause a lock loss now when trying to lock the arm with high bandwidth using POY.

  12146   Thu Jun 2 16:35:44 2016 KojiUpdateSUS wire standoffs update

Gap of the prism from the mirror

Sag: s = R(1-Cos[ArcSin[d/2/R]])

- Mirror curvature sag for 2mm prism (R=37.75mm): s=13um

- Minimum gap: 20um => s=33um => R=15mm

- Nominal gap: 35um => s=48um => R=10mm

- Maximum gap: 50um => s=63um => R=8mm


The second figure shows somewhat realistic arrangement of the pieces

Attachment 1: sapphire_stand_off_v3.pdf
sapphire_stand_off_v3.pdf
Attachment 2: arrangement-v3.pdf
arrangement-v3.pdf
  12145   Wed Jun 1 16:28:28 2016 ericqUpdateElectronicsCommon board Op amp input offsets

I used a Eurocard extension board to peek at the inputs and outputs of each of the gain-ladder AD829s on input B of the CM board in the +31dB configuration with the input terminated. (i.e with the following stages active in this order: +16dB, +8dB, +4dB, +2dB, +1dB).

The voltages I observed imply that the +8dB stage has an input voltage offset of -2mV, whereas all the other positive gain stages show around +-0.5mV. This could explain the shift observed at the +15->+16 transition. (However, since both input channels show a jump here, maybe its something more systemic about the board...)

In any case, it should be simple enough to swap out a new AD829 in place of U9B and see if it improves things, before getting too deep into the muck. (In principle, the AD829 has offset nulling pins, but I'm not sure how to do it in a non-hacky way since the board doesn't have any pads for it.)

  12144   Wed Jun 1 15:01:56 2016 SteveUpdateSUS wire standoffs update

There are some issues with 5 mm sapphire prism Atm5. It will cause  interference between one of the prisms and the Side OSEM.

Here are some drawings to see the issues with larger wire standoff.

The 2 mm prism will work.with a 1 mm longer dumbell.

Quotes requested from http://photomachining.com/laser-micromachining-photomachining-contact.html and http://www.optocity.com/ 

 

 

Attachment 1: 06011603.PDF
06011603.PDF
Attachment 2: 06011602.PDF
06011602.PDF
Attachment 3: sapphire_stand_off_v2.pdf
sapphire_stand_off_v2.pdf
Attachment 4: side_OSEM_ETMY.jpg
side_OSEM_ETMY.jpg
Attachment 5: 5p2wsf.png
5p2wsf.png
  12143   Wed Jun 1 11:19:14 2016 VarunUpdateGeneralUpdate of work till now

Completed:

Wrote and tested a code for AGC using cavity transmission signal and length error signal.

Wrote and tested a code for frequency shifting (downconversion) using mixing and LPF

Wrote a code for whitening using FFT.

Altium working on cit40m iMac

Plans:

Writing codes for Frequency warping and whitening in time domain.

Implement AGC and frequency shifting on the real time control system.

Calculate requirements for Anti-aliasing filter.

  12142   Wed Jun 1 09:06:38 2016 SteveUpdateLSCNew stands for TransMon/Oplev QPDs

Machined from I-beam 6061 T6 Aluminum 5" x 0.5 x 3.25

Quote:

As we realized during the EX table switch, the transmitted beam height from the arm is not exactly 4" relative to the endtable, it is more like 4.75" at the X-end (yet to be investigated at the Y-end). As a result, the present configuration involves the steering optics immediately before the Oplev and TransMon QPDs sending the beam downwards at about 5 degrees. Although this isn't an extremely large angle, we would like to have things more level. For this purpose, Steve has ordered some Aluminium I-beams (1/2 " thick) which we can cut to size as we require. The idea is to have the QPD enclosures mounted on these beams and then clamped to the table. One concern was electrical isolation - but Steve thinks Delrin washers between the QPD enclosure and the mount will suffice. We will move ahead with getting these machined once I investigate the situation at the Y end as well.. The I beams should be here sometime next week...

Atm2, version 2 "pdstand" will allow you to clamp from any direction ( Koji was right )

Attachment 1: pdIb.PDF
pdIb.PDF pdIb.PDF
Attachment 2: pdstand.pdf
pdstand.pdf
  12141   Tue May 31 16:52:58 2016 SteveUpdatesafetyNONO

Please do not place anything on the top of the cabinets that is not tied down. It will end up on our head in an earth quake.

 

Attachment 1: nono.jpg
nono.jpg
  12140   Mon May 30 18:19:50 2016 JohannesUpdateCDSASS medm screen update

I noticed that the TRY button in the ASS main screen was linking to LSC_TRX instead of LSC_TRY. Gautam fixed it.

  12139   Fri May 27 11:54:22 2016 VarunUpdateGeneralPackage delivery

A package labelled 'UPS Ground' has arrived.

-Varun

  12138   Fri May 27 02:52:53 2016 ericqUpdateLSCRestoring high BW single arm control

I've been futzing with the common mode servo, trying to engage the AO path with POY for high bandwidth control of a single arm lock. I'm able to pull in the crossover and get a nice loop shape, but keep getting tripped up by the offset glitches from the CM board gain steps, so can't get much more than a 1kHz UGF.

As yutaro measured, these can be especially nasty at the major carrier transitions (i.e. something like 0111->1000). This happens at the +15->+16dB input gain step; the offset step is ~200x larger than the in-loop error signal RMS, so obviously there is no hope of keeping the loop engaged when recieving this kind of kick. Neither of the CM board inputs are immune from this, as I have empirically discovered. I can turn down the initial input gain to try and avoid this step occuring anywhere in the sequence, but then the SNR at high frequencies get terrible and I inject all kinds of crud into the mode cleaner, making the PC drive furious.

I think we're able to escape this when locking the full IFO because the voltages coming out of REFL11 are so much larger than the puny POY signals so the input-referred glitches aren't as bad. I think in the past, we used AS55 with a misaligned ITMX for this kind of single arm thing, which probably gives better SNR, but the whole point of this is to keep the X arm aligned and lock it to the Y-arm stabilized PSL. 

  12137   Thu May 26 18:10:48 2016 VarunUpdateGeneralSURF 2016

Wrote and tested a function for downconversion. It contains a mixer with a sinusoidal input for modulation with the desired frequency and a 2nd order butterworth low pass filter to remove the higher frequency-shifted part of the modulated signal. I have tested this with input of 2kHz giving a good output of 200 Hz on the speaker. Codes are uploaded on github, will update the real time document tomorrow.

 

-Varun

Quote:

Edited the AGC to include overlapping frames yesterday. forgot to put an elog on it!

Quote:

Tested the AGC today with LSC cavity transmission signal and error signal. Not in real time still.

Key to attachments:

cav_tr-eps-converted-to.pdf: LSC cavity transmission signal input

cav_tr_out-eps-converted-to.pdf: LSC cavity transmission signal, output of the AGC.

 

 

Attachment 1: input.png
input.png
Attachment 2: output.png
output.png
  12136   Wed May 25 14:29:31 2016 VarunUpdateGeneralSURF 2016

Edited the AGC to include overlapping frames yesterday. forgot to put an elog on it!

Quote:

Tested the AGC today with LSC cavity transmission signal and error signal. Not in real time still.

Key to attachments:

cav_tr-eps-converted-to.pdf: LSC cavity transmission signal input

cav_tr_out-eps-converted-to.pdf: LSC cavity transmission signal, output of the AGC.

 

  12135   Wed May 25 14:21:29 2016 Max IsiUpdateGeneralSummary page configuration

I have modified the c1summary.ini and c1lsc.ini configuration files slightly to avoid overloading the system and remove the errors that were preventing plots from being updated after certain time in the day.

The changes made are the following:
1- all high-resolution spectra from the Summary and LSC tabs are now computed for each state (X-arm locked, Y-arm locked, IFO locked, all);
2- I've removed MICH, PRCL & SRCL from the summary spectrum (those can still be found in the LSC tab);
3- I've split LSC into two subtabs.

The reason for these changes is that having high resolution (raw channels, 16kHz) spectra for multiple (>3) channels on a single tab requires a *lot* of memory to process. As a result, those jobs were failing in a way that blocked the queue, so even other "healthy" tabs could not be updated.

My changes, reflected from May 25 on, should hopefully fix this. As always, feel free to re organize the ini files to make the pages more useful to you, but keep in mind that we cannot support multiple high resolution spectra on a single tab, as explained above.

  12134   Wed May 25 11:51:40 2016 SteveUpdatesafetySURF 2016 safety
Quote:

Hello, I am Varun Kelkar. I will be working at the 40m lab as a SURF student this summer with Eric Quintero on Audio processing for real time control system signals. This week I will mostly be working on implementing basic DSP C-code offline. Currently I am trying to write a code for noise whitening.

-Varun

Varun has received 40m specific basic safety training today.

  12133   Wed May 25 08:32:55 2016 SteveUpdateSUSlocal EQ 3.5m

Local EQ 3.5 mag  at 2:28 UTC May 24, 2016 Rancho Cucamonga, Ca.....no damage

 

Attachment 1: 3.5Cucam.png
3.5Cucam.png
Attachment 2: local3.5cucam.png
local3.5cucam.png
  12132   Wed May 25 02:54:09 2016 ericqUpdateGeneralOdds and ends

WFS locking point seemed degraded; I hand aligned and reset the WFS offsets as usual.

ITMX oplev recentered. While doing so, I noticed an ETMX excursion rear its head for the first time in a long while :crying

There was no active length control on ETMX, only OSEM damping + oplevs. Afterwards, its still moving around with only local damping on. I'm leaving the oplevs off for now.

  12131   Tue May 24 23:17:37 2016 ericqUpdateCOCFinesse modelling - mode overlap scans

I think you should use the current actual PRC & SRC cavity lengths as measured, as it would be simplest to simply replace the folding mirror optics without changing the macroscopic lengths / optic positions. (EDIT: Gautam rightly points out that we have to move things around regardless, since our current lengths include propagation through the folding mirror subtrates)

Moreover, the recycling cavity lengths you posted are not the right "ideal" lengths to use, as they do not account for the complex reflectivities of the sidebands off of the arm cavities (I have made this mistake myself). See this 40m wiki page for details.

In short, given our current modulation frequency, the ideal lengths to use would be:

  • Ideal arm length of 37.795 m
  • Ideal PRC length of 6.753 m
  • Ideal SRC length of 5.399 m

These are the lengths that the recycling cavity optics were positioned for (though we did not achieve them perfectly). If you do a finer PRC/SRC length scan around the DRFPMI resonance of your model, you would presumably see some undesired sideband splitting. 

  12130   Tue May 24 22:49:02 2016 gautamUpdateCOCFinesse modelling - mode overlap scans

Summary:

Having played around with a toy finesse model, I went about setting up a model in which the RC folding mirrors are not flipped. I then repeated the low-level tests detailed in the earlier elog, after which I ran a few spatial mode overlap analyses, the results of which are presented here. It remains to do a stability analysis.

Overview of model parameters (more details to follow):

  • PRC length = 6.7727m (chosen using l_{PRC} = (N+\frac{1}{2})\frac{c}{2f_1}, N=0 - I adjusted the position of the PRM to realize this length in the model, while leaving all the other vertex optics in the same positions as in elog 9590
  • SRC length = 5.4182 (chosen using l_{SRC} = M\frac{c}{2f_2} but not l_{SRC} = N\frac{c}{2f_1}, M and N being integers, for M=2 - as above, I adjusted the position of the SRM to realize this in the model, while leaving all other vertex optics in the same positions as in elog 9590. It remains to be verified if it is physically possible to realize these dimensions in vacuum without any beam clipping etc but I think it should be possible seeing as the PRM and SRM had to be moved by less than 2cm from their current positions..
  • For the losses, I used the most recent numbers we have where applicable, and put in generic 25ppm loss for all the folding mirrors/BS/AR surfaces of arm cavity mirrors/PRM/SRM. Arm round trip loss was equally distributed between ITMs and ETMs
  • Arm lengths used: L_X = 37.79m, L_Y = 37.81m
  • To set the "tunings" of the various mirrors, I played around with a few configurations to see where the various fields resonated - it turns out that for PRM, ITMX, ITMY, ETMX and ETMY, the "phase" in the .kat file can be set as 0. while that for the SRM can be set as 90. In the full L1/H1 interferometer .kat files, these are tuned even further to the (tenth?!) decimal place, but I think these values suffice for out purposes.

Results (general note: positive RoC in these plots mean a concave surface as seen by the beam):

  • Attachments #1, #2 and #3 reproduce the low-level tests performed earlier for this updated model - i.e. I look at the arm transmission with no PRM/SRM, circulating PRC power with no ETMs, and circulating SRC power with no ETMs. Everything looks consistent here... In Attachment #2, there is no legend, but the (almost overlapping) red and green lines are meant to denote the +f1 and +f2 sidebands.
  • Attachments #4 and #5 are a summary of the mode-overlap scans for the PRC and SRC. What I did was to vary the radius of curvature of the RC mirrors (finesse only allows you to vary Rcx and Rcy, so I varied both simultaneously) and calculate the mode overlap between the appropriate pairs of cavities (e.g. PRX and XARM) in the tangential and saggital planes. The take-away here is that there is ~5% mode-mismatch going from an RoC of 1000m to 300m. I've also indicated the sag corresponding to a given RoC - these are pretty tiny, I wonder if it is possible to realize a sag of 1um? I suppose it is given that I've regularly seen specs of surface roughness of lambda/10?
  • Attachment #6 shows the PRC gain (calculated as T_PRC * (transmitted arm power with PRM / transmitted arm power without PRM) as a function of the RoC of PR2 and PR3. As a sanity check, I repeated this calculation with lossless HR surfaces (but with nominal 25ppm losses for AR surfaces of ITMs, and BS etc), shown in Attachment #7. I think these make sense too...
  • Attachment #8 - in order to investigate possible mode mismatch between the arm modes due to different radii of curvature of the ETMs, I kept the ETMY RoC fixed at 57.6m and varied the ETMY RoC between 50m and 70m (here, I've plotted the mode matching efficiency as a function of the RoC of the ETM in the X and Y directions separately - the mode overlap is computed as \frac{1}{\sqrt{2}}(x^2 + y^2) where x and y denote the overlap in the tangential and saggital planes respectively. It would seem that we only lose at most a couple of percent even if the RoCs are mismatched by up to 10m...
  • Attachment #9 - .kat file and the various pykat scripts used to generate these plots...

Next step is to carry out a stability analysis...

Attachment 1: armTransmission.pdf
armTransmission.pdf
Attachment 2: prcFSR.pdf
prcFSR.pdf
Attachment 3: srcTransmission.pdf
srcTransmission.pdf
Attachment 4: modeMatchPRX.pdf
modeMatchPRX.pdf
Attachment 5: modeMatchSRX.pdf
modeMatchSRX.pdf
Attachment 6: PRCgainScan.pdf
PRCgainScan.pdf
Attachment 7: PRCgainLossless.pdf
PRCgainLossless.pdf
Attachment 8: armModeMatchScan.pdf
armModeMatchScan.pdf
Attachment 9: Finesse_files.zip
  12129   Tue May 24 17:55:17 2016 VarunUpdateElectronicsUsing Altium

Contacted Charles regarding use of Altium. Got to know that Altium is installed on cit40m iMac in Win7 on VirtualBox. Had to update Virtualbox to get it working. Altium now works for sometime, but then fails, saying that it is unlicensed.

  12128   Tue May 24 10:21:36 2016 ericqSummarySUSITMX Oplev loops

I did a quick measurement of the ITMX oplev loops, both pitch and yaw have about the same upper UGF as previous measurements with the previous laser; about 4 Hz. 

  12127   Mon May 23 17:47:51 2016 VarunUpdateGeneralSURF 2016

Tested the AGC today with LSC cavity transmission signal and error signal. Not in real time still.

Key to attachments:

cav_tr-eps-converted-to.pdf: LSC cavity transmission signal input

cav_tr_out-eps-converted-to.pdf: LSC cavity transmission signal, output of the AGC.

Attachment 1: cav_tr-eps-converted-to.pdf
cav_tr-eps-converted-to.pdf
Attachment 2: cav_tr_out-eps-converted-to.pdf
cav_tr_out-eps-converted-to.pdf
Attachment 3: err-eps-converted-to.pdf
err-eps-converted-to.pdf
Attachment 4: err_out-eps-converted-to.pdf
err_out-eps-converted-to.pdf
  12126   Mon May 23 15:51:32 2016 steveSummarySUSoplev laser summary updated

 

Quote:

 

Quote:

 

Quote:

 

                  2005              ALL oplev servos use Coherent DIODE LASERS # 31-0425-000, 670 nm, 1 mW

    Sep. 28, 2006              optical lever noise budget with DC readout in 40m,  LIGO- T060234-00-R, Reinecke & Rana

    May  22, 2007              BS, SRM & PRM  He Ne 1103P takes over from diode

    May  29, 2007              low RIN He Ne JDSU 1103P selected, 5 purchased sn: T8078254, T8078256, T8078257, T8078258 & T8077178 in Sep. 2007

    Nov  30, 2007               Uniphase 1103P divergence measured

    Nov. 30, 2007               ETMX old Uniphase 1103P  from 2002 dies: .............., running time not known......~3-5 years?

    May 19, 2008               ETMY old Uniphase 1103P from 1999 dies;.....................running time not known.....~    ?

    Oct.  2, 2008                ITMX & ITMY are still diodes, meaning others are converted to 1103P earlier

 

                     JDSU 1103P were replaced as follows:

   May 11, 2011                ETMX replaced, life time 1,258 days  or 3.4 years

   May 13, 2014               ETMX , LT 1,098 days or 3 y

   May 22, 2012               ETMY,  LT 1,464 days or  4 y

   Oct.  5, 2011                BS & PRM, LT 4 years,  laser in place at 1,037 days or 2.8 y

   Sep. 13, 2011               ITMY  old 1103P &    SRM    diode laser replaced by 1125P  ..........old He life time is not known, 1125P in place 1,059 days or 2.9 y

   June 26, 2013              ITMX 622 days or 1.7 y    note: we changed because of beam quality.........................laser in place 420 days or 1.2 y

 

  Sep. 27, 2013               purchased 3 JDSU 1103P lasers, sn: P893516, P893518, P893519 ......2 spares ( also 2 spares of 1125P of 5 mW & larger body )

 

      May  13, 2014             ETMX,  .............laser in place 90 d

      May  22, 2012             ETMY, 

     Oct.  7,  2013             ETMY,  LT  503 d  or  1.4 y............bad beam quality ?

     Aug. 8,  2014              ETMY,  .............laser in place   425 days  or  1.2 y

 

      Sept. 5, 2014              new 1103P, sn P893516  installed at SP table for aLIGO oplev use qualification

     

           May 23, 2016             ITMX dead laser sn P845648 replaced after 1062 days [2.9 yrs] by 1103P, sn P859884, with output output  2.6 mW, nicely round beam quality at 15 meters.

Attachment 1: oplSum.png
oplSum.png
  12125   Mon May 23 10:55:49 2016 steveSummarySUSITMX oplev laser replaced

 

      May 23, 2016             ITMX dead He/Ne laser sn P845648 replaced after 1062 days [2.9 yrs] by 1103P, sn P859884, with output  2.6 mW, nicely round beam quality at 15 meters.

                                                                                                                                                    Power just before viewport 1 mW,  returning light on qpd 154 microW =  7,500 counts

 

Attachment 1: ITMXoplev.png
ITMXoplev.png
  12124   Fri May 20 17:36:06 2016 gautamUpdateLSCNew stands for TransMon/Oplev QPDs

As we realized during the EX table switch, the transmitted beam height from the arm is not exactly 4" relative to the endtable, it is more like 4.75" at the X-end (yet to be investigated at the Y-end). As a result, the present configuration involves the steering optics immediately before the Oplev and TransMon QPDs sending the beam downwards at about 5 degrees. Although this isn't an extremely large angle, we would like to have things more level. For this purpose, Steve has ordered some Aluminium I-beams (1/2 " thick) which we can cut to size as we require. The idea is to have the QPD enclosures mounted on these beams and then clamped to the table. One concern was electrical isolation - but Steve thinks Delrin washers between the QPD enclosure and the mount will suffice. We will move ahead with getting these machined once I investigate the situation at the Y end as well.. The I beams should be here sometime next week...

  12123   Fri May 20 00:06:19 2016 VarunUpdateGeneralSURF 2016

I have written a basic version of AGC, and have done some tests with a data file. will do tests on whitening and agc today. Also, today I have to go to the SSN office. Hence will be late.

 

-Varun

Quote:

Finished writing the code on whitening. I have to still test it. uploaded on github noise cancellation repo. @eric could you give me some data of noise power spectral density for testing the code?

-Varun

Quote:

Hello, I am Varun Kelkar. I will be working at the 40m lab as a SURF student this summer with Eric Quintero on Audio processing for real time control system signals. This week I will mostly be working on implementing basic DSP C-code offline. Currently I am trying to write a code for noise whitening.

-Varun

 

 

  12122   Thu May 19 16:29:20 2016 SteveUpdateendtable upgradeOptical layout almost complete

 

 

Attachment 1: ETMX_4x3_closed.jpg
ETMX_4x3_closed.jpg
Attachment 2: sealedETMXenclosure.jpg
sealedETMXenclosure.jpg
  12121   Wed May 18 17:42:52 2016 VarunUpdateGeneralSURF 2016

Finished writing the code on whitening. I have to still test it. uploaded on github noise cancellation repo. @eric could you give me some data of noise power spectral density for testing the code?

-Varun

Quote:

Hello, I am Varun Kelkar. I will be working at the 40m lab as a SURF student this summer with Eric Quintero on Audio processing for real time control system signals. This week I will mostly be working on implementing basic DSP C-code offline. Currently I am trying to write a code for noise whitening.

-Varun

 

  12120   Wed May 18 01:10:22 2016 gautamUpdateCOCFinesse modelling

I've been working on putting together a Finesse model for the current 40m configuration. The idea was to see if I could reproduce a model that is in agreement with what we have been seeing during the recent DRFPMI locks. With Antonio and EricQs help, I've been making slow progress in my forays into Finesse and pyKat. Here is a summary of what I have so far.

  • Arm lengths were taken from some recent measurements done by yutaro and me 
  • Recycling cavity lengths were taken from Gabriele's elog 9590 - it is likely that the lengths I used have errors ~1cm - more on this later. Furthermore, I've tried to incorporate the flipped RC folding mirrors - the point being to see if I can recover, for example, a power recycling gain of ~7 which is what was observed for the recent DRFPMI locks.
  • I used Yutaro's most recent arm loss numbers, and distributed it equally between ITM and ETM for modeling purposes. 
  • For all other optics, I assumed a generic loss number of 25ppm for each surface

Having put together the .kat file (code attached, but this is probably useless, the new model with RC folding mirrors the right way will be what is relevant), I was able to recover a power recycling gain of ~7.5. The arm transmission at full lock also matches the expected value (125*80uW ~ 10mW) based on a recent measurement I did while putting the X endtable together. I also tuned the arm losses to see (qualitatively) that the power recycling gain tracked this curve by Yutaro. EricQ suggested I do a few more checks:

  1. Set PRM reflectivity to 0, scan ETMs and look at the transmission - attachment #1 suggests the linewidth is as we expect 
  2. Set ETM reflectivity to 0, scan PRM - attachment #2 suggests a Finesse of ~60  for the PRC which sounds about right
  3. Set ETM reflectivity to 0, scan SRM and verify that only the 55 MHz sidebands resonate - Attachment #3

Conclusion: It doesn't look like I've done anything crazy. So unless anyone thinks there are any further checks I should do on this "toy" model, I will start putting together the "correct" model - using RC folding mirrors that are oriented the right way, and using the "ideal" RC cavity lengths as detailed on this wiki page. The plan of action then is

  • Evaluating the mode-matching integrals between the PRC and the arm cavities as a function of the radius of curvature of PR2 and PR3
  • Same as above for the SRC
  • PRC gain as a function of RoC of folding mirrors
  • Mode overlap between the modes from the two arm cavities as a function of the RoC of the two ETMs (actually I guess we can fix RoC of ETMy and just vary RoC of ETMx).

Sidenote to self: It would be nice to consolidate the most recent cavity length measurements in one place sometime...

Attachment 1: arms.pdf
arms.pdf
Attachment 2: PRC.pdf
PRC.pdf
Attachment 3: SRC.pdf
SRC.pdf
Attachment 4: Finesse_model.zip
  12119   Tue May 17 14:46:51 2016 SteveUpdateVACRGA scan at day 595

We have good RGA scan now. There was no scan for 3 months.

Attachment 1: RGAscan595d.png
RGAscan595d.png
Attachment 2: pd78-560Hz-d600.png
pd78-560Hz-d600.png
  12118   Tue May 17 05:50:43 2016 Varun KelkarUpdateGeneralSURF 2016

Hello, I am Varun Kelkar. I will be working at the 40m lab as a SURF student this summer with Eric Quintero on Audio processing for real time control system signals. This week I will mostly be working on implementing basic DSP C-code offline. Currently I am trying to write a code for noise whitening.

-Varun

  12117   Sun May 15 19:48:08 2016 SteveUpdateVACrun out of N2

3-4 hrs ago we run out of nitrogen. We are back to Vacuum Normal

 

 

Attachment 1: noN2.png
noN2.png
  12116   Thu May 12 14:29:58 2016 gautamUpdateVACRGA back up and running

It looks like the hardware reset did the trick. Previously, I had just tried ssh-ing into c0rga and rebooting it. At the time, however, Steve and I noticed that the various LEDs on the RGA unit weren't on, as they are supposed to be in the nominal operating state. Today, Steve reported that all LEDs except the RS232 one were on today, so I just tried following the steps in this elog again, looks like things are back up and running. I'm attaching a plot of the scan generated using plotrgascan MATLAB script, it looks comparable to the plot in elog 11697, which if I remember right, was acceptable.

Unless there is some reason we want to keep this c0rga machine, I will recommission one of the spare Raspberry Pis lying around to interface with the RGA scanner when I get the time...

Quote:
Quote:

Our last RGA scan is from February 14, 2016  We had a power outage on the 15th

Gautom has not succeded  reseting it. The old c0rga computer looks dead. Q may resurrect it, if he can?

The c0rga computer was off, I turned it on via front panel button. After running RGAset.py, RGAlogger.py seems to run. However, there are error messages in the output of the plotrgascan MATLAB script; evidiently there are some negative/bogus values in the output. 

I'll look into it more tomorrow.

This is a cold scan.

Attachment 1: RGAscan_12May2016.png
RGAscan_12May2016.png
  12115   Wed May 11 16:39:01 2016 ericqUpdateVACc0rga alive, output wonky
Quote:

Our last RGA scan is from February 14, 2016  We had a power outage on the 15th

Gautom has not succeded  reseting it. The old c0rga computer looks dead. Q may resurrect it, if he can?

The c0rga computer was off, I turned it on via front panel button. After running RGAset.py, RGAlogger.py seems to run. However, there are error messages in the output of the plotrgascan MATLAB script; evidiently there are some negative/bogus values in the output. 

I'll look into it more tomorrow.

  12114   Tue May 10 03:44:59 2016 ericqUpdateLSCRelocked

ALSX noise is solidly within past acceptable performance levels. The DRFPMI was locked on four out of six attempts. 

Some housekeeping was done:

  • PMC aligned
  • Static alignment voltages of X end PZT mirrors offloaded by turning mount screws
  • Rough comissioning of AUX X dither alignment
  • Locking scripts reverted to AUX X Innolight voltage/temperature sign convention

The recombination of the QPD signals to common / differential is imperfect, and limited how well we could keep the interferometer aligned, since the QPD at X has changed. This needs some daytime work. 

Some sensing matrix measurements were made, to be meditated upon for how to 1F the DRMI.

Other to-dos:

  • Bandpass + notch combo for green refl PDs
  • SRCL, and to a lesser extant, MICH feedforward subtraction (See DARM vs. other length DOF coherence plot below)
  • Fiber couple AUX X light
  • Make IFO work good


As an aside, Gautam and I noticed numerous green beams coming from inside the vacuum system onto the PSL table. They exist only when green is locked to the arms. Some of them come out at very non-level angles and shine in many places. This doesn't make me feel very happy; I suppose we've been living with it for some time. 

Attachment 1: 2016-05-10_DARMcoherence.pdf
2016-05-10_DARMcoherence.pdf
  12113   Sun May 8 08:39:21 2016 ranaUpdateLSCGreen PDH demod lowpass

Indeed. This is why the LSC PDs have a 2f notch in addition to the 1f resonance. In recent versions, we also put a 2f notch in the feedback of the preamp which comes after the diode but before the mixer. The overall 1f to 2f ratio that we get is in the 50-60 dB region. I don't think we have to go that far with this thing; having a double LC already seems like it should be pretty good, or we could have a single LC bandpass with a 2f notch all in one Pomona box.

  12112   Sat May 7 09:40:40 2016 ericqUpdateLSCGreen PDH demod lowpass

As I was looking at filter designs, it seemed difficult to get 40dB of supression at 2F with a bandpass without going to a pretty high order, which would mean a fair number of lossy inductors.

I'll keep working on it. Maybe we don't need 40dB...

  12111   Fri May 6 19:08:52 2016 ranaUpdateLSCGreen PDH demod lowpass

Seems weird to design a PD lowpass with a corner at the modulation frequency. Recall what our strategy is with the other photodetectors we use for PDH servos: bandpass, not low-pass, and the band has to be wide enough to not effect the phase of the servo.

  12110   Fri May 6 16:42:12 2016 ericqUpdateLSCGreen PDH demod lowpass

I've build the filter, and it seems to have the desired TF shape.

I also re-purposed the 70k lowass to a ~120k lowpass by changing the 68nF caps to 22nF caps, since we still want some post-mixer rolloff. 

However, putting the ELPF in the chain caused some weird shapes in the OLG. I still need to get to the bottom of it. However, just with the post-mixer LPF modification, here's what the OLG looks like:

As Rana surmises, we definitely still add a boost and maintain a 10k UGF. I still need to look into the state of the remote boost....

  12109   Thu May 5 21:28:44 2016 gautamUpdateendtable upgradeInnolight PZT capacitance

I suggested in an earlier elog that after the repair of the NPRO, the PZT capacitance may have changed dramatically. This seems unlikely - I measured the PZT capacitance with the BK Precision LCR meter and found it to be 2.62 nF, which is in excellent agreement with the numbers from elogs 3640 and 4354 - but this makes me wonder how the old setup ever worked. If the PZT capacitance were indeed that value, then for the Pomona box design in elog 4354, and assuming the PM at ~216kHz which was the old modulation frequency was ~30rad/V as suggested by the data in this elog, we would have had a modulation depth of 0.75 if the Function Generator were set to output a Signal at 2Vpp (2Vpp * 0.5 * 0.05 * 30rad/V = 1.5rad pp)! Am I missing something here?

Instead of using an attenuator, we could instead change the capacitor in the pomona box from 47pF mica to 5pF mica to realize a modulation depth of ~0.2 at the new modulation frequency of 231.25 kHz. In any case, as elog 4354 suggests, the phase introduced by this high-pass filter is non-zero at the modulation frequency, so we may also want to install an all-pass filter which will allow us to control the demodulation phase. This should be easy enough to implement with an Op27 and passive components we have in hand...

 

  12108   Thu May 5 14:05:01 2016 ranaUpdateendtable upgradeALS status update

All seems very fishy. Its not good to put attenuators and filters in nilly-willy.

  1. Once the post-PD bandpass has been designed and constructed, you should be able to use whatever PD gain setting gives you the best SNR. There's no need to use more PD gain than necessary; it just reduces the PD bandwidth. What is the input referred current noise of the PD at the different gain settings?
  2. The open loop mixer output *should* be very large. It should be reduced to mV only when the loop is closed.
  3. The better way to estimate the modulation depth is to lock the arm on red as usual and then scan the EX laser and look at the green transmission. The FSR is 3.7 MHz, so the SBs should show up well in a narrow scan around the carrier.
  4. I guess its going to be tough to impedance match the splitter box to the NPRO PZT, since its impedance is all over the place at 200-300 kHz, but you could put a 50 Ohm in-line terminator in there somewhere?
  5. The Bode plot seems to indicate that we could easily get a 10 kHz UGF and then switch on a Boost. Is the remote Boost switch disabled or always ON? I am suspicious of the plot and think that the coarse trace is probably missing some sharp resonances which will sneakily bite you.
  12107   Thu May 5 14:03:52 2016 ericqUpdateLSCFurther Aux X PDH tweaks

This morning I poked around with the green layout a bit. I found that the iris immediately preceding the viewport was clipping the ingoing green beam too much, opening it up allowed for better coupling to the arm. I also tweaked the positions of the mode matching lenses and did some alignment, and have since been able to achieve GTRX values of around 0.5.

I also removed the 20db attenuator after the mixer, and turned the servo gain way down and was able to lock easily. I then adjusted the gain while measuring the CLG, and set it where the maximum gain peaking was 6dB, which worked out to be a UGF of around 8kHz. On the input monitor, the PDH horn-to-horn voltage going into the VGA is 2.44V, which shouldn't saturate the G=4 preamp stage of the AD8336, which seems ok.

The ALS sensitivity is now approaching the good nominal state:

There remains some things to be done, including comprehensive dumping of all beams at the end table (especially the reflections off of the viewport) and the new filters to replace the current post-mixer LPF, but things look pretty good.

Attachment 1: 2016-05-05_newals.pdf
2016-05-05_newals.pdf
  12106   Thu May 5 04:05:03 2016 ericqUpdateLSCAux X PDH checks

We took an OLG measurement of the green PDH loop. It seems consistent with past measurements. I've added a trace for the the post-mixer lowpass, to show its contribution to the phase loss. (EDIT: updated with measured LPF TF)

I used this measured OLG and the datasheet laser PZT conversion factor to calibrate the control signal monitor into the AUX laser frequency noise, it looks consistent with the frequency noise measured via the PSL PLL (300 Hz/rtHz @ 100Hz). Above a few tens of kHz, the control signal measurement is all analyzer noise floor, due to the fourth order 70kHz lowpass after the mixer (the peaks change height significantly depending on the analyzer input range, so I don't think they're on the laser). Gautam will follow up with more detailed measurements of both the error and control signals as he noisebudgets, this was just intended as a quick consistency check.

  12105   Thu May 5 03:05:37 2016 gautamUpdateendtable upgradeALS status update

[ericQ, gautam]

Today we spent some time looking into the PDH situation at the X end. A summary of our findings.

  1. There is something that I don't understand with regards to the modulation signal being sent to the laser PZT via the sum+HPF pomona box - it used to be that with 2Vpp signal from the function generator, we got ~5mVpp signal at the PZT, which with the old specs resulted in a modulation of ~0.12rad. Now, however, I found that there was a need to place a 20dB attenuator after the splitter from the function generator in order to realize a modulation depth of ~0.25 (which is what we aim for, measured by locking to the TEM00 modes of the carrier and sidebands and comparing the ratio of powers). It could be that the PZT capacitance has changed dramatically after the repair. Nevertheless, I still cant reconcile the numbers. We measured the transfer function from the LO input of the pomona box to the output with the PZT connected, and figure there should be ~70dB of attentuation (with the 20dB additional attenuator in place). But this means 1Vpp*0.0003*70rad/V = 0.02rad which is an order of magnitude away from what the ratio of powers suggest. Maybe the measurement technique was not valid. In any case, this setup appears to work, and I'm also able to send +7dBm to the mixer which is what it wants (function generator output is 3Vpp).
  2. In addition to the above, I found that the demodulated error signal had a peak-to-peak of a few volts. But the PDH servo is designed to have tens of mV at the input. Hence, it was necessary to turn down the gain of the REFL PD to 10dB and add a 20dB attenuator between mixer output and servo input.
  3. While Johannes and I were investigating this earlier in the afternoon, we found that the waveform going to the laser PZT was weirdly distorted (still kind of sinusoidal in shape, but more rounded, I will put up a picture shortly). This may not be the biggest problem, but perhaps there is a better way to pipe the LO signal to the PZT and mixer than what is currently done.
  4. We then looked at loop transfer function and spectrum of the control signal. Plots to follow. They look okay.
  5. I measured the green power coming onto the PSL table. It is ~400uW. After optimizing alignment, the green transmission is ~0.4 according to whatever old normalization we are using.
  6. We then recovered the X green beatnote and looked at the ALS noise spectrum. Beatnote amplitude at the beat PD is ~ -27dBm. The coherence in the region of a few hundred Hz suggests that some improvements can be made to the PDH situation (the gain of the PDH servo is maxed out at the X end at the moment...). But the bottom line is this is probably good enough to get back to locking...
Attachment 1: ALS_noiseSpec_5May2016_2.pdf
ALS_noiseSpec_5May2016_2.pdf
Attachment 2: Coherence_5May2016.pdf
Coherence_5May2016.pdf
Attachment 3: image.jpeg
image.jpeg
  12104   Mon May 2 19:14:18 2016 gautamUpdateendtable upgradeOptical layout almost complete

With Steve's help, I installed the Oplev earlier today. I adjusted the positions of the two lenses until I deemed the spot size on the QPD satisfactory by eye. As a quick check, I verified using the DTT template that the UGF is ~5Hz for both pitch and yaw. There is ~300uW of power incident on the QPD (out of ~2mW from the HeNe). In terms of ADC counts, this is ~13,000 counts which is about what we had prior to taking the endtable apart. There are a couple of spots from reflections off the black glass plate in the vacuum chamber, but in general, I think the overall setup is acceptable.

This completes the bulk of the optical layout. The only bits remaining are to couple the IR into the fiber and to install a power monitoring PD. Pictures to follow shortly. 

Now that the layout is complete, it remains to optimize various things. My immediate plan is to do the following:

  1. Maximize green transmission by tweaking alignment. I should also do a quick check using mirror specs to see that the measured transmitted green power compares favourably to what is expected.
  2. Check the green PDH loop transfer function at the X end - this will allow me to set the gain on the uPDH box systematically.
  3. Re-establish green beats, check noise performance.
  4. There are possibly multiple beam dumps that have to be installed. For now, I've made sure that no high power IR beams are incident on the enclosure. But there are a couple of red and green beams that have to be accounted for.

I will also need to upload the layout drawing to reflect the layout finally implemented.


Not directly related:

The ETMx oplev servo is now on. I then wanted to see if I could lock both arms to IR. I've managed to do this successfully - BUT I think there is something wrong with the X arm dither alignment servo. By manually tweaking the alignment sliders on the IFOalign MEDM screen, I can get the IR transmission up to ~0.95. But when I run the dither, it drives the transmission back down to ~0.6, where it plateaus. I will need to investigate further. 

 

GV Edit: There was some confusion while aligning the Oplev input beam as to how the wedge of the ETM is oriented. We believe the wedge is horizontal, but its orientation (i.e. thicker side on the right or left?) was still ambiguous. I've made a roughly-to-scale sketch (attachment #1) of what I think is the correct orientation - which turns out to be in the opposite sense of the schematic pinned up in the office area.. Does this make sense? Is there some schematic/drawing where the wedge orientation is explicitly indicated? My search of the elog/wiki did not yield any..

Attachment 1: ETMX_wedge.pdf
ETMX_wedge.pdf
  12103   Mon May 2 17:11:55 2016 ranaUpdateCOCRC folding mirrors

Antonio/Gautam are now developing a more up to date Finesse model of our recycling cavities to see what we can have there before our power recycling gain or cavity geometric stability is compromised. Expect that we will here a progress report on the model on Wednesday.

Some thoughts:

  1. RC folding mirrors need to be dichroic to allow green beams to get out.
  2. We should look at the specs Jamie used to get the RC folding mirrors last time and figure out what went wrong / what specs to change.
  3. T_1064 < 100 ppm. Hopefully < 50 ppm.
  4. On the AR side, we mainly want low AR for green, but nothing special for 1064, since that's taken care of by the HR.
  5. How much should we wedge these things?
  6. Should the wedge be horizontal?
  7. Can we get someone in Downs to update the optical layout?
  8. What microroughness do we need?
  9. The mirrors must be flat, with the  500 m < RoC < 100 km. Part of the Finesse modeling is to figure out what happens if the RoC is in the 300 - 1000 m range. Better stability?
  12102   Mon May 2 17:06:58 2016 ranaSummaryCOCG&H optics to Fullerton/HWS for anneal testing

Steve sent 4 of our 1" diameter G&H HR mirrors to Josh Smith at Fullerton for scatter testing. Attached photo is our total stock before sending.

Attachment 1: 20160427_182305.jpg
20160427_182305.jpg
  12101   Fri Apr 29 16:13:36 2016 ericqUpdateLSCGreen PDH demod lowpass

We can get as much, if not more, attenuation of the 1F line in the mixer output that we get from the post-mixer LPF from using the following passive filter between the PD and mixer RF input:

There should still be some kind of LPF after the mixer, but I haven't yet determined what it should be; this will determine how much phase the PDH loop wins. At most, this should win around 25 degrees at 10kHz.


The filter was designed by referencing the "Handbook of Filter Synthesis" by Zverev, looking for an elliptic filter for matched source and load impedences, 40dB min attenuation in the stopband, a stopband frequency that starts at twice the corner frequency, and minimizing the VSWR between the PD and filter in the passband.

In terms of the tables in the book, this means: n=5, rho=2%, theta=30deg, K**2 = 1.0. The dimensionless component values were scaled by the corner frequency of 200kHz, and reference impedence of 50 Ohm. (The corner is a little lower than the real modulation frequency, since the nonzero resistance of the inductors pushes the frequency up a bit)

The ideal capactior values do not correspond to things we have in hand, so I checked our stock and chose the closest value to each one.Unsurprisingly, due to these component substitutions, and the fact that the coilcraft inductors have a resistance of about 7 Ohms, the predicted TF of the realizable filter does not match the design filter exactly. However, the predicition still looks like it will meet the requirement of 40dB of supression of the 2F line in the PD signal. (Since we have tunable inductors, I've used the ideal inductor values in generating the TF. In practice I'll inspect the TF while I tune them)

  Desired Realizable
C1 8.28 nF 10 nF
C2 1.39 nF 1.5 nF
C3 19.6 nF 22 nF
C4 4.22 nF 4.7 nF
C5 6.08 nF 6.8 nF
L2 43.1 nH 32-48 nH + 7 Ohm
L4 34.4 nH 32-48 nH + 7 Ohm

[In this TF plot, I've multiplied the real response by 2 to account for the voltage division that occurs with ideal 50 Ohm impedance matching, to make 0dB the reference for proper matching]

The filter's phase delay at the modulation frequency is just about 180, which as a time delay of 5usec works out to 9 degrees of phase loss at 10kHz in the PDH loop. According to some old measurements, the current LPF costs something like 35 degrees at 10k, so this wins at most around 25 degrees, depedent on what LPF we put after the mixer.

LISO source both traces is attached!

Attachment 3: elp_liso.zip
  12100   Fri Apr 29 16:05:23 2016 gautamUpdateendtable upgradeCleaning ETMX vacuum dirty window

After a second round of F.C. application, I think the window is clean enough and there are no residual F.C. pieces anywhere near the central parts of the window (indeed I think we got most of it off). So I am going to go ahead and install the Oplev. 

Quote:

It looks very promising.

 

 

Attachment 1: IMG_0755.JPG
IMG_0755.JPG
  12099   Fri Apr 29 00:55:46 2016 gautamUpdateendtable upgradegreen PDH locked to Xarm

Using the modulation frequency suggested here, I hooked up the PDH setup at the X-end and succeeded in locking the green to the X arm. I then rotated the HWP after the green Faraday to maximize TRX output, which after a cursory alignment optimization is ~0.2 (I believe we were used to seeing ~0.3 before the end laser went wonky). Obviously much optimization/characterization remains to be done. But for tonight, I am closing the PSL and EX laser shutters and applying first contact to the window once more courtesy more PEEK from Koji's lab in W Bridge. Once this is taken care of, I can install the Oplev tomorrow, and then set about optimizing various things in a systematic way.. MC autolocker has also been disabled...

Side note: for the IR Transmon QPD, we'd like a post that is ~0.75" taller given the difference in beam height from the arm cavity and on the endtable. I will put together a drawing for Steve tomorrow..

ELOG V3.1.3-